Points A and B are situated along the extended axis of a 2cm long bar magnet at a distance x and 2x cm respectively. From the pole nearer to the points, the ratio of the magnetic field of A and B will be?

Grade:12th Pass

2 Answers

Arun
4 years ago
B proportional to 1/ x cube

B1/B2 = (2x) cube / x cube = 8:1
 
Hope it helps
 
Regards
Arun (askIITians forum expert)
Khimraj
4 years ago

B1/B2 = (x1/x2)3 = (x/2x)3 = 1:8
Hope it clears..............................................................

B1:B2 = 27:1